OpenStudy (anonymous):

How do I do this? Find the ratio of X toy Y X/5=2/3=5/Y

OpenStudy (anonymous):

find x then find y \[x/5=2/3\] \[5/y=2/3\]

OpenStudy (anonymous):

How do you do that? Would you cross multiply or divide?

OpenStudy (anonymous):

yes cross mutliply

OpenStudy (anonymous):

So for X I would get 15/3?

OpenStudy (anonymous):

Is that right? @Jonask

OpenStudy (anonymous):

X=5*2/3=10/3 Y=5*3/2=15/2